OpenStudy (julius):

My second time on this site. Please help. Does anyone know if this is a typo. Its OCW Scholar Multivariable calculus Session 62. The PDF title is "Gradient Fields". Page 2 midpage!. F = del ln times square root of x^2 + y^2. Shouldn't that be 1/2 of del ln times square root of x^2 + Y^2? When I differentiate with respect to x and then y there is a 2 in the denominator so don't you need to take it out with the 1/2? Or am I really missing the boat on this? Thank you guys and gals. Julius

OpenStudy (phi):

I think there is a typo (missing minus sign) where they take the partials (but it does not matter much). Specifically \[ \frac{\partial}{\partial x } \left( \frac{y}{x^2+y^2}\right) \\ =\frac{\partial}{\partial x } y(x^2+y^2)^{-1} = -y(x^2+y^2)^{-2} 2 x\\ =\frac{-2xy}{(x^2+y^2)^2} \] similarly for the other component (there is a missing minus sign)

OpenStudy (phi):

as for the gradient of \( \ln\sqrt{x^2+y^2} \) we use \[ d \ln u = \frac{1}{u} \ du\] the first component is \[ \frac{\partial}{\partial x} (x^2+y^2)^\frac{1}{2} \\ = \frac{1}{(x^2+y^2)^{\frac{1}{2}}} \cdot \frac{1}{2} (x^2+y^2)^ {-\frac{1}{2}}\cdot 2x\\ = \frac{x}{x^2+y^2} \] similarly for the other component

OpenStudy (julius):

Thank you so much!!! you are so right and so kind! off to the next section! Have a good day